# LoKI: Low-damage Knowledge Implanting of Large Language Models

### v1

@@ 3.2.1 Layer-Balanced Strategy
3.3 Implanting
Building upon the selected knowledge output nodes $\mathcal{S}=\bigcup_{l=1}^{L}\mathcal{S}_{l}$ from Section 3.2, we implement parameter-efficient fine-tuning through strategic decomposition of FFN layers. For each layer $l$’s down-projection matrix $W_{\mathrm{down}}^{(l)}$, we partition the parameters into two complementary subspaces:
@@ 3.3 Implanting
This decomposition retains the mathematical formulation of the original layer while restricting parameter updates to the chosen subspaces. Additionally, it transforms LoKI training into a module-wise process, enabling easy integration with existing training pipelines. Notably, this implementation allows for the incorporation of low-rank decomposition techniques [hu2022lora, DBLP:conf/icml/LiuWY0WCC24, DBLP:conf/iclr/ZhangCBH0CZ23] into LoKI. To verify the feasibility of this combination, we conducted an experiment in Section 4.2. For clarity, we refer to this implementation as LoKI Linear, and readers can find its PyTorch code in Appendix C.
4 Experiments
In this section, we detail the experiments conducted to evaluate the effectiveness of the LoKI. We focus on two tasks: adapting models on the LB Reranker Dataset and the ToolACE Function-Calling Dataset. These publicly available datasets were specifically selected due to their demonstrated practical utility in real-world applications and their ability to simulate realistic PEFT demand scenarios. Each experiment assesses the capability of LoKI to mitigate catastrophic forgetting while acquiring task-specific performance. Notably, while our experiments involved only two model types, the proposed method is readily applicable to other model architectures.
@@ 4 Experiments
    LB Reranker Dataset [lb_reranker]: This multilingual dataset consists of 2.28 million query-text pairs annotated with fine-grained relevance scores on a 1-7 scale, designed for training NLP-based retrieval models. It has an official full parameter fine-tuning model based on Qwen2.5-0.5B-Instruct [qwen_qwen25_2025]; we utilized the same base model on this dataset. The experiment setup for this task can be found in Appendix D.
@@
    ToolACE Function-Calling Dataset [liu_toolace_2024]: This dataset contains 26,507 distinct APIs across various domains, designed to enhance the function-calling capabilities of LLMs. This richly diverse and rigorously validated dataset underpins state-of-the-art performance in function-calling benchmarks [berkeley-function-calling-leaderboard], even for 8B-parameter models. Consistent with the official model, we fine-tuned Llama3.1-8B-Instruct [grattafiori_llama_2024] on this dataset. The details of the experiment setup can be found in Appendix E.
For both tasks, to systematically evaluate CF during fine-tuning, we employ six benchmarks to assess model capabilities across diverse domains: TriviaQA [DBLP:conf/acl/JoshiCWZ17] (world knowledge), GSM8K [cobbe2021gsm8k] (mathematical reasoning), HellaSwag [DBLP:conf/acl/ZellersHBFC19] and WinoGrande [DBLP:conf/aaai/SakaguchiBBC20] (commonsense understanding), HumanEval [chen2021evaluating] (code generation), and IFEval [zhou_instruction-following_2023] (instruction following). The information regarding our evaluation of these benchmarks is in Appendix F.
4.1 Task 1: LB Reranker Dataset
Table 1: Comparison with the full parameter fine-tuning model on BEIR.
LoKI($q{=}5$) -3.1 -5.9 -3.1 -3.5 -2.2 -4.0 -1.8 -3.6
LoKI($q{=}10$) -2.3 -3.8 -2.3 -2.1 -0.3 -2.2 -0.5 -2.0
LoKI($q{=}20$) 0.2 -2.5 0.2 -1.5 0.7 -1.4 0.8 -1.5
LoKI($q{=}30$) 1.0 -1.0 1.0 -0.5 2.1 -0.3 2.5 -0.5
Qwen2.5 -78.1 -77.7 -78.1 -68.3 -78.7 -71.8 -77.5 -65.4
Table shows the percentage difference in performance compared to the full parameter fine-tuning model across standard retrieval metrics. Bold represents the top performance score in each column; underline represents the runner-up. Qwen2.5 represents Qwen2.5-0.5B-Instruct
The BEIR benchmark [thakur_beir_2021] is a standardized evaluation framework for testing the performance of information retrieval models across a wide range of diverse datasets and tasks. Experimental results on the BEIR benchmark (Table 14) demonstrate that our LoKI method achieves competitive performance compared to full parameter fine-tuning while requiring substantially fewer trainable parameters. Specifically, LoKI with $q{=}30$ surpasses the full-parameter LB-Reranker baseline by an average of 0.54% across all evaluation metrics. Notably, it demonstrates absolute improvements in enhanced precision for the top-ranked retrieval results.
The progressive performance improvement with increasing trainable parameters (from $q{=}5$ to $q{=}30$) reveals a clear parameter-performance tradeoff. Even with only $q{=}20$, LoKI already achieves positive performance gains in four metrics.
Table 2: Benchmark scores of models fine-tuned on LB Reranker Dataset.
Model #Params TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(%, ↓)
Qwen2.5 NaN 24.37 39.65 30.98 44.20 26.83 33.39 NaN
LoKI($q{=}5$) 5.1M 20.53 38.36 30.52 40.09 24.39 34.12 6.12
LoKI($q{=}10$) 10.4M 20.53 37.98 21.53 47.04 24.39 33.39 8.86
LoKI($q{=}20$) 20.9M 19.77 36.77 33.16 47.28 27.44 33.27 1.70
LoKI($q{=}30$) 31.3M 20.10 37.60 38.39 43.96 26.83 32.25 0.46
LB-Reranker 494M 4.65 2.65 0.00 0.00 0.00 23.20 84.13
LB-Reranker refers to the full parameter fine-tuning model released by the dataset provider. Avg denotes the average performance degradation percentage of each indicator compared to the original model indicator.
As shown in Table 2, LoKI models of all configurations exhibit significantly less performance degradation across all metrics compared to the full parameter fine-tuned model. Interestingly, when $q{=}30$, LoKI not only achieved the best averaged performance on the BEIR benchmark, but also showed the least degradation on general tasks. We believe this situation is related to the learning rate settings (refer to Appendix D for further discussion).
4.2 Task 2: ToolACE Function-Calling Dataset
Model Overall Acc (%, ↑) Single Turn Acc Multi Turn (%, ↑) Hallucination Measurement
Non-Live (%, ↑) Live (%, ↑) Relevance (%, ↑) Irrelevance (%, ↓)
ToolACE 58.32 87.56 76.10 7.62 83.33 88.05
LoKI($q{=}10$) 56.76 80.81 70.46 16.50 83.33 82.21
LoKI($q{=}20$) 58.16 81.02 73.26 17.75 83.33 85.73
LoKI($q{=}30$) 58.93 82.71 75.70 16.25 83.33 87.65
LoKI*($q{=}30$) 57.16 81.96 71.66 15.75 83.33 84.01
Table 3: Performance comparison on Berkeley Function Calling Leaderboard V3. ToolACE denotes the official model trained using LoRA. * indicates that we incorporated LoRA during training in this model.
Model #Params TriviaQA GSM8K Hellaswag WinoGrande HumanEval IFEval Avg(%, ↓)
Llama3.1 NaN 65.77 84.46 73.85 62.98 68.29 79.76 NaN
LoKI($q{=}10$) 188M 65.85 84.99 75.14 61.33 68.90 77.54 0.34
LoKI($q{=}20$) 376M 65.60 84.31 73.60 60.62 67.68 78.18 0.93
LoKI($q{=}30$) 563M 65.49 84.61 71.16 61.09 70.73 77.94 1.23
LoKI*($q{=}30$) 16M 64.30 83.47 70.78 62.51 70.12 78.21 1.26
ToolACE 42M 64.63 79.53 46.81 42.78 60.37 72.76 16.11
Table 4: Benchmark scores of models fine-tuned on ToolACE Function-Calling Dataset. Llama3.1 denotes Llama3.1-8B-Instruct.
Table 3 demonstrates that models trained using LoKI exhibit performance comparable to ToolACE, which was trained using LoRA on all modules. Specifically, LoKI($q{=}30$) achieves the highest overall accuracy (58.93%), surpassing the official ToolACE baseline. Notably, LoKI($q{=}20$) exhibits the strongest multi-turn reasoning capability with a 17.75% success rate, significantly outperforming ToolACE’s mere 7.62%. All LoKI variants reduce indicators of Irrelevance, which may suggest the ability of LoKI to alleviate the model’s hallucination after fine-tuning.
Table 4 further reveals LoKI’s unique capability in preserving foundational model competencies. Despite ToolACE’s 16.11% average performance drop across six benchmarks, LoKI($q{=}30$) maintains near-original capabilities with only 1.23% degradation, particularly evident in complex reasoning tasks like GSM8K (84.61% vs ToolACE’s 79.53%) and programming challenges (70.73% HumanEval accuracy vs 60.37%). This preservation consistency holds across all model scales.
Additionally, we experimented to investigate the potential of combining LoRA with LoKI (LoKI*($q{=}30$)). As anticipated, the number of trainable parameters in the model significantly decreased when integrated with LoRA, showing a reduction of 97.16% compared to LoKI($q{=}30$). Importantly, this combination did not visibly compromise LoKI’s ability to resist catastrophic forgetting. Although its performance on specialized tasks after fine-tuning is not as high as that of other models, we believe that this combination holds significant promise for future optimization.
5 Ablation Studies
In this section, we evaluate (1) the precision of KVA in quantifying knowledge vectors’ contributions to general task performance, and (2) the effectiveness of the Layer-Balanced Strategy in protecting model general task performance.
5.1 Effect of Knowledge Vector Attribution
We use Llama3.1-8B-Instruct to validate if KVA can effectively attribute the contribution of knowledge vectors to general tasks. We compare the base model with two suppression strategies (i.e., zeroing the corresponding weights):
  • S-H/L: Suppresses the top $q$knowledge vectors with the highest/lowest attribution scores (expected to significantly/minimally degrade performance if KVA correctly identifies critical vectors)
Table 5: Performance of two suppression strategies on benchmarks when $q{=}1$.
Model TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(%, ↓)
Llama3.1 65.77 84.46 73.85 62.98 68.29 79.76 NaN
S-H($q{=}1$) 62.32 15.16 23.54 60.93 43.29 59.73 36.73
S-L($q{=}1$) 62.05 57.85 73.84 60.3 54.88 74.15 11.35
As presented in Table 5, when the top 1% of high-contribution vectors are suppressed, there is a significant performance decline, with an average degradation of 36.73%. In contrast, the performance gap between the two suppression strategies is 25.38%, which indicates significant differences and supports the accuracy of KVA.
5.2 Effect of Layer-Balanced Strategy
Next, we test the necessity of the Layer-Balanced Strategy by fine-tuning Qwen2.5-0.5B-Instruct on the LB Rerank Dataset, using two globally imbalanced schemes. We compare these results with models trained using LoKI on evaluation benchmarks.
  • G-H/L: Globally set the top $q$knowledge vectors across all layers with the highest/lowest attribution scores as trainable parameters.
As shown in Table 6, both imbalanced strategies underperform compared to LoKI: G-H shows a 39.04% average performance gap, while G-L exhibits a 30.48% average gap. The performance gaps observed in both imbalanced strategies, despite utilizing the same quota as LoKI, demonstrate the necessity of the Layer-Balanced Strategy in LoKI. We also provide the performance on the BEIR benchmark for these two models in Appendix G.
Table 6: Performance comparison of different methods.
Model TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(%, ↓)
LoKI($q{=}10$) 20.53 37.98 21.53 47.04 24.39 33.39 8.86
G-H($q{=}10$) 11.04 29.34 10.23 17.68 25.00 26.79 39.04
G-L($q{=}10$) 9.83 30.93 8.48 43.09 24.39 27.72 30.48
6 Related Work
Catastrophic Forgetting. The challenge of catastrophic forgetting has been extensively studied in neural networks [DBLP:conf/nips/French93, kemker2018measuring], with recent works specifically analyzing its manifestation in LLMs [li_revisiting_2024, DBLP:conf/iclr/KothaSR24]. The existing methods to mitigate CF in LLMs include rehearsal-based approaches [DBLP:conf/nips/dAutumeRKY19, DBLP:conf/nips/Lopez-PazR17, DBLP:conf/iclr/ChaudhryRRE19] and regularization strategies [DBLP:journals/pami/LiH18a, DBLP:journals/corr/KirkpatrickPRVD16, deng_flattening_2021]. However, the understanding of the model’s native knowledge storage mechanism is rarely mentioned in the process of resisting CF.
Parameter-Efficient Fine-Tuning. By updating only a small subset of parameters, PEFT reduces tuning overhead: Adapters [DBLP:conf/icml/HoulsbyGJMLGAG19, DBLP:conf/iclr/HeZMBN22] insert bottleneck modules in each transformer layer; LoRA [hu2022lora] injects low-rank matrices into attention/FFN weights; prefix tuning [DBLP:conf/acl/LiL20, DBLP:journals/corr/abs-2110-07602] adds learnable "virtual tokens." Recent studies have been dedicated to improving the issue of catastrophic forgetting in PEFT methods [DBLP:conf/coling/LengX25, DBLP:journals/corr/abs-2503-02659, DBLP:conf/emnlp/WangCGXBZZGH23, DBLP:journals/corr/abs-2503-13575], but these primarily focus on the forgetting problem that arises between multiple subtasks rather than protecting the original capabilities of the model.
Knowledge Locating and Editing. In recent years, there has been a surge in techniques for modifying the knowledge of LLMs [zhang_comprehensive_2024]. For instance, ROME [meng_locating_2022] modifies the FFN weight vectors to revise one factual association, and KN [DBLP:conf/acl/DaiDHSCW22] controls the activation level of knowledge neurons to regulate the expression of relevant facts. Geva et al. [geva_transformer_2021-1, geva_dissecting_2023, geva_transformer_2022-1] conducted extensive research on how transformers intrinsically store knowledge, providing theoretical guidance for knowledge localization and editing. However, a gap remains in integrating such mechanisms into efficient fine-tuning pipelines that preserve the model’s broad pre-trained capabilities.
7 Conclusions
We present LoKI, a parameter-efficient fine-tuning framework that achieves balanced adaptation between downstream task performance and preservation of pre-trained knowledge in large language models. Our key insight stems from systematically analyzing the hierarchical knowledge storage mechanism in transformer FFN layers and developing a layer-balanced parameter selection strategy guided by integrated gradient attribution. Through experiments on retrieval and tool-use tasks, we demonstrate that LoKI achieves competitive task adaptation while significantly reducing catastrophic forgetting compared to full parameter fine-tuning and LoRA. By connecting insights from mechanistic interpretability research to practical fine-tuning goals, our study emphasizes the promising potential of this interdisciplinary field.

### v2

@@ Introduction
To empirically assess the effectiveness of the Layer-Balanced Strategy, we perform an ablation study comparing it with imbalanced baselines that naïvely select knowledge vectors with the highest or lowest global attribution scores as trainable parameters. Experimental details and results are provided in the Ablation Studies section.
Implanting
Building upon the selected knowledge output nodes $\mathcal{S}=\bigcup_{l=1}^{L}\mathcal{S}_{l}$, we implement parameter-efficient fine-tuning through strategic decomposition of FFN layers. For each layer $l$’s down-projection matrix $W_{\mathrm{down}}^{(l)}$, we partition the parameters into two complementary subspaces:
@@
This decomposition retains the mathematical formulation of the original layer while restricting parameter updates to the chosen subspaces. In addition, it transforms LoKI training into a module-wise process, allowing easy integration with existing training pipelines. For clarity, we refer to this implementation as LoKI Linear, and the readers can find its PyTorch code in Appendix C.
Incorporating LoRA
It is worth noting that this implementation allows for the incorporation of low-rank decomposition techniques (hu2022lora; DBLP:conf/icml/LiuWY0WCC24; DBLP:conf/iclr/ZhangCBH0CZ23) into LoKI. Specifically, the trainable weights are parameterized as:
$W_{\mathcal{S}_{l}}=W_{\mathcal{S}_{l}}^{(0)}+\Delta W_{\mathcal{S}_{l}},\quad\Delta W_{\mathcal{S}_{l}}=A_{l}B_{l},$ (11)
where $W_{\mathcal{S}_{l}}^{(0)}$ denotes the frozen base weights, and $A_{l}\in\mathbb{R}^{|\mathcal{S}_{l}|\times r}$, $B_{l}\in\mathbb{R}^{r\times d_{\mathrm{ff}}}$ are learnable low-rank matrices with rank $r\ll\min(|\mathcal{S}_{l}|,d_{\mathrm{ff}})$. To verify the feasibility of this integration, we experimented with the Experiment section.
Experiments
To evaluate our proposed method, we focus on adapting models to the following two datasets:
@@
    ToolACE Function-Calling Dataset (liu_toolace_2024): This dataset contains 26,507 distinct APIs across various domains, designed to enhance the function-calling capabilities of LLMs. Consistent with the official model released by the dataset provider, we fine-tuned Llama3.1-8B-Instruct on this dataset. The details of the experiment setup can be found in Appendix D.
@@
    LB Reranker Dataset (lb_reranker): This multilingual dataset consists of 2.28 million query-text pairs annotated with fine-grained relevance scores on a 1-7 scale, designed for training NLP-based retrieval models. It has an official full parameter fine-tuning model based on Qwen2.5-0.5B-Instruct (qwen_qwen25_2025); we utilized the same base model on this dataset. The experiment setup for this task can be found in Appendix E.
These publicly available datasets were specifically selected due to their demonstrated practical utility in real-world applications and their ability to simulate realistic fine-tuning demand scenarios. Each experiment assesses the capability of LoKI to resist the CF phenomenon while acquiring task-specific performance. Notably, while our experiments involved only two model types, the proposed method is readily applicable to other model architectures.
Model
Overall Acc
Single Turn Acc
Multi Turn
Hallucination Measurement
Non-Live(%, ↑)
Live(%, ↑)
Relevance(%, ↑)
Irrelevance(%, ↓)
@@
DoRA(r=16r{=}16) 58.90 82.04 75.61 16.00 83.33 88.92
PiSSA(r=16r{=}16) 53.97 80.56 72.68 5.62 61.11 92.19
LoKI(q=10q{=}10) 56.76 80.81 70.46 16.50 83.33 82.21
LoKI(q=20q{=}20) 58.16 81.02 73.26 17.75 83.33 85.73
LoKI(q=30q{=}30) 58.93 82.71 75.70 16.25 83.33 87.65
LoKI*(q=30,r=32q{=}30,r{=}32) 57.16 81.96 71.66 15.75 83.33 84.01
Table 1: Performance comparison on the Berkeley Function Calling Leaderboard V3. ToolACE refers to the official model trained using LoRA. $r$ denotes the rank. Models marked with $\dagger$ indicate that the corresponding performance metrics are sourced directly from the official BFCL documentation. An asterisk (*) denotes models where LoRA was applied during training. Bold represents the top performance score in each column; underline represents the runner-up.
Model #Params TriviaQA GSM8K Hellaswag WinoGrande HumanEval IFEval Avg(↓)
Llama3.1 (untuned) 65.77 84.46 73.85 62.98 68.29 79.76
ToolACE(r=16r{=}16) 42M 64.63 79.53 46.81 42.78 60.37 72.76 16.11%
DoRA(r=16r{=}16) 43M 63.97 82.64 70.78 57.85 65.24 73.47 4.92%
PiSSA(r=16r{=}16) 42M 52.50 51.40 37.98 13.18 15.85 55.95 48.93%
LoKI(q=10q{=}10) 188M 65.85 84.99 75.14 61.33 68.90 77.54 0.34%
LoKI(q=20q{=}20) 376M 65.60 84.31 73.60 60.62 67.68 78.18 0.93%
LoKI(q=30q{=}30) 563M 65.49 84.61 71.16 61.09 70.73 77.94 1.23%
LoKI*(q=30,r=32q{=}30,r{=}32) 16M 64.30 83.47 70.78 62.51 70.12 78.21 1.26%
Table 2: Benchmark scores of models fine-tuned on ToolACE Function-Calling Dataset. Llama3.1 denotes Llama3.1-8B-Instruct. Avg denotes the average performance degradation percentage of each indicator compared to the original model indicator. We complete all the benchmarks on OpenCompass and report their results directly.
Model MAP@1 (%) Recall@1 (%) NDCG@1 (%) P@1 (%)
Qwen2.5 -78.1 -78.1 -78.7 -77.5
DoRA(r=8r{=}8) -8.7 -8.7 -11.2 -10.4
PiSSA(r=8r{=}8) -9.6 -9.6 -9.7 -10.2
CorDA(r=8r{=}8) -3.3 -3.1 -2.6 -3.3
LoKI(q=5q{=}5) -3.1 -3.1 -2.2 -1.8
LoKI(q=10q{=}10) -2.3 -2.3 -0.3 -0.5
LoKI(q=20q{=}20) +0.2 +0.2 +0.7 +0.8
LoKI(q=30q{=}30) +1.0 +1.0 +2.1 +2.5
Table 3: Retrieval performance on BEIR benchmark. Results show the percentage difference relative to the full parameter fine-tuning model across standard retrieval metrics, where positive values indicate superior performance. CorDA uses the KPM setting on the NQ-Open dataset.
Measuring Catastrophic Forgetting
To systematically evaluate CF during fine-tuning, we assess the model’s retained general capabilities across six diverse benchmarks: TriviaQA (DBLP:conf/acl/JoshiCWZ17) (world knowledge), GSM8K (cobbe2021gsm8k) (mathematical reasoning), HellaSwag (DBLP:conf/acl/ZellersHBFC19) and WinoGrande (DBLP:conf/aaai/SakaguchiBBC20) (commonsense understanding), HumanEval (chen2021evaluating) (code generation), and IFEval (zhou_instruction-following_2023) (instruction following).
We define the average forgetting score across all benchmarks as: $\text{Avg}=\frac{100}{N}\sum_{i=1}^{N}\frac{S_{i}^{\text{o}}-S_{i}^{\text{t}}}{S_{i}^{\text{o}}},$ where $S_{i}^{\text{o}}$ and $S_{i}^{\text{t}}$ denote the performance of the original and fine-tuned models on the $i$-th benchmark, respectively, and $N$ is the total number of benchmarks. A higher value of Avg indicates a greater degree of forgetting. Further details regarding the evaluation of these benchmarks can be found in Appendix F.
Task 1: ToolACE Function-Calling Dataset
Model #Params TriviaQA GSM8K HellaSwag WinoGrande HumanEval IFEval Avg(↓)
Qwen2.5 (untuned) 24.37 39.65 30.98 44.20 26.83 33.39
LB-Reranker 494M 4.65 2.65 0.00 0.00 0.00 23.20 84.13%
DoRA(r=8r{=}8) 4M 19.30 34.01 22.21 44.70 25.22 31.67 12.23%
PiSSA(r=8r{=}8) 4M 11.32 6.44 18.54 31.97 6.71 28.85 48.95%
CorDA(r=8r{=}8) 4M 4.64 2.43 0.00 0.00 0.00 23.20 84.22%
LoKI(q=5q{=}5) 5.1M 20.53 38.36 30.52 40.09 24.39 34.12 6.12%
LoKI(q=10q{=}10) 10.4M 20.53 37.98 21.53 47.04 24.39 33.39 8.86%
LoKI(q=20q{=}20) 20.9M 19.77 36.77 33.16 47.28 27.44 33.27 1.70%
LoKI(q=30q{=}30) 31.3M 20.10 37.60 38.39 43.96 26.83 32.25 0.46%
Table 4: Benchmark scores of models fine-tuned on the LB Reranker Dataset. LB-Reranker refers to the full parameter fine-tuning model released by the dataset provider. Qwen2.5 represents the Qwen2.5-0.5B-Instruct.
Based on the training results on this dataset, we first compare LoKI with existing SOTA methods in terms of downstream task adaptation. We evaluate fine-tuned models’ performance on the Berkeley Function Calling Leaderboard V3, comparing LoKI against LoRA (model provided by the dataset authors), DoRA (DBLP:conf/icml/LiuWY0WCC24), and PiSSA (DBLP:conf/nips/MengWZ24). As shown in Table 2, LoKI achieves the highest overall accuracy when q=30q{=}30. Under the q=20q{=}20 setting, LoKI exhibits the strongest multi-turn reasoning capability, with a success rate of 17.75%. Notably, all LoKI variants consistently reduce the Irrelevance metric, suggesting that LoKI may help mitigate hallucination effects introduced during fine-tuning. Table 2 further reveals LoKI’s unique capability in preserving foundational model competencies. Compared to DoRA, LoKI (q=30q{=}30) not only outperforms in fine-tuned performance, but also reduces the average performance degradation percentage by 75% across six evaluation benchmarks. Relative to the original pretrained model, the degradation is as small as 1.23%. Notably, as the hyperparameter $q$ increases from 10 to 30—despite a substantial growth in the number of trainable parameters—the rate of performance degradation slows down. Additionally, we experimented to investigate the potential of combining LoRA with LoKI (LoKI*(q=30q{=}30)). As anticipated, the number of trainable parameters in the model significantly decreased when integrated with LoRA, showing a reduction of 97.16% compared to LoKI(q=30q{=}30). Importantly, this combination did not visibly compromise LoKI’s ability to resist catastrophic forgetting. we believe that this combination holds significant promise with further optimization of training settings.
Task 2: LB Reranker Dataset
[…]
Ablation Studies
Model Avg(↓)
S-H 36.73%
S-L 11.35%
Model Avg(↓)
LoKI 8.86%
G-H 39.04%
G-L 30.48%
Table 5: (a) Performance of two suppression strategies on benchmarks when q=1q{=}1. (b) Performance comparison of different methods when q=10q{=}10. Detailed results are provided in Appendix G.
In this section, we evaluate (a) the effect of KVA in quantifying knowledge vectors’ contributions to general task performance, and (b) the necessity of the Layer-Balanced Strategy in LoKI.
Validation of KVA
We use Llama3.1-8B-Instruct to validate if KVA can effectively attribute the contribution of knowledge vectors to general tasks. We compare two suppression strategies (i.e., zeroing the corresponding weights) with the base model using the same benchmarks in the Experiments section:
  • S-H/L: Suppresses the top qq% knowledge vectors that appear most frequently with the highest/lowest attribution scores (expected to significantly/minimally degrade performance if KVA correctly identifies critical vectors).
As presented in Table 5(a), when the top 1% of high-contribution vectors are suppressed, there is a significant performance decline, with an average degradation of 36.73%. In contrast, the performance gap between the two suppression strategies is 25.38%, which indicates significant differences and supports the accuracy of KVA.
Validation of Layer-Balanced Strategy
Next, we test the necessity of the Layer-Balanced Strategy by fine-tuning Qwen2.5-0.5B-Instruct on the LB Rerank Dataset, using two globally imbalanced schemes. Equally, we compare these results with models trained using LoKI on the 6 benchmarks mentioned above.
  • G-H/L: Globally set the top qq% knowledge vectors across all layers that appear most frequently with the highest/lowest attribution scores as trainable parameters.
As shown in Table 5(b), both imbalanced strategies fall short of LoKI, despite using the same quota, underscoring the importance of the Layer-Balanced Strategy. Additional BEIR benchmark results are included in Appendix G.
Conclusions
We present LoKI, a parameter-efficient fine-tuning framework that achieves balanced adaptation between downstream task performance and preservation of pre-trained knowledge in large language models. Our key insight stems from systematically analyzing the hierarchical knowledge storage mechanism in transformers and developing a layer-balanced parameter selection strategy guided by integrated gradient attribution. Through experiments on retrieval and tool-use tasks, we demonstrate that LoKI achieves competitive task adaptation while significantly reducing catastrophic forgetting compared to full parameter fine-tuning and prevalent PEFT methods. Our experimental results demonstrate that integrating insights from mechanistic interpretability research with fine-tuning objectives is effective, highlighting the potential of this interdisciplinary direction.
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grants 52202496, 52442218, and U2433216; The Key Research and Development Project of Nantong City, China (Special Project for Prospective Technology Innovation, No. GZ2024001); and the Key Laboratory of Target Cognition and Application Technology (2023-CXPT-LC-005).